A supply plane needs to drop a package of food to scientists working on a glacier in Greenland. The plane fies 140 m above the glacier at a speed of 200 m/s How far short of the target should it drop the package? Express your answer using three significant figures with the appropriate units You may want to review (Pages 91-95) For help with math skills, you may want to review Mathematical Expressions Involving Squares Value Units For general problem-solving tips and strategies for this topic, you may want to view a Video Tutor Solution of Dock jumping
Answer #1

In y direction:

a = -9.8 m/s^2

d = -140 m

vi = 0 m/s

use:

d = vi*t + 0.5*a*t^2

-140 = 0 + 0.5*(-9.8)*t^2

4.9*t^2 = 140

t = 5.345 s

In horizontal direction,

distance moved by package = v*time

= 200 m/s * 5.345 s

= 1070 m

So, it should be dropped 1070 m or 1.07 km ahead

Answer: 1.07 km